OpenStudy (anonymous):

a mixture contains 12 litters of waters and 3 litters of alcohol. If 6 more litters of water is added, what part of the mixture is alcohol?

OpenStudy (anonymous):

are you working out the percentage of alcohol?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

21 litres total, 3 of which are alcohol. 3/21 = 1/7. Or 3:18 = 1:6 alcohol-water ratio.

OpenStudy (anonymous):

where did you get 21 its 12

OpenStudy (anonymous):

i need another answer

OpenStudy (anonymous):

There's 12 litres of water. Add another 6 litres. That makes 18 litres. Plus 3 litres of alcohol makes 21.

OpenStudy (anonymous):

they 1/7 but how

OpenStudy (anonymous):

Because there's 21 litres total: 12 + 3 + 6 = 21. Of that, 3 litres are alcohol which is 3/21 which simplifies to 1/7 (divide 3 and 21 by 3).

OpenStudy (anonymous):

why did you add the alcohol 3

OpenStudy (anonymous):

Because it's part of the mixture. As a whole, it is 21 litres, including the 3 litres of alcohol. If it were separate, it wouldn't be a part at all.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

i try the total of water:12 litres alcohol:3 litres 12/3=4 litres add 6 litres=6-4 =2 so,myb the answer is 2 litres